OpenStudy (anonymous):

solve the equation for exact solutions. sin^-1x+2tan^1x=pi

OpenStudy (anonymous):

do you mean sin^-1x+2tan^-1x=pi?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the solution of 1 can be found via guess & check; not sure how to find it algebraically

OpenStudy (anonymous):

sorry, im still trying to solve it

OpenStudy (anonymous):

is this meant to be solved using calculus?

OpenStudy (anonymous):

no we are doing this in trigonometry...dealing with arcsin, arccos, and arctan, my professor likes to switch around problems so its hard to undestand them

OpenStudy (alekos):

yes , it can be solved. what have you got?

OpenStudy (anonymous):

well what i did is someow move arcsin x to pi...but i dont think that is right

OpenStudy (alekos):

doing it in stages we get 1/sinx + 2/tanx = pi 1/sinx + 2cosx/sinx = pi (1 + 2cosx)/sinx = pi 1 + 2cosx = pi*sinx pi*sinx - 2cosx = 1 my mistake, it doesnt come down to a simple solution let me think about it

OpenStudy (anonymous):

ok no problem

OpenStudy (alekos):

this can't be solved algebraically, i think you'll need to do it graphically

OpenStudy (alekos):

been thinking about this one.... for y = asinx +bcosx we can find y = Asin(x+c) where a = Acosc and b = Asinc in this case a=pi and b=-2, so we can solve for A and c then we solve for Asin(x+c) =1 and we get x = arcsin(1/A) - c give this a try and see what you get